OpenStudy (anonymous):

PLEASE HELP FAN AND MEDAL IN RETURN I NEED HELP I NEED TO GET THIS DONE PLEASE HELP!!

OpenStudy (icecreamyummy):

whats the question

OpenStudy (anonymous):

This project may be completed individually or collaboratively. The town of West Mathington needs help planning some roads that will connect parts of their fair city. The town of West Mathington is laid out so that North is the positive y-axis, and East is the positive x-axis. The roads will follow the paths of graphs created from linear, quadratic, and exponential functions. These samples provided will give you an idea of what a final map could look like.

OpenStudy (anonymous):

1. West Mathington’s most urgent need is a parabolic freeway. Create your own upward opening quadratic function, f(x), which has two real zeros. Prove that f(x) has two real zeros. 2. Two on-ramps need to be placed on the parabolic freeway. Decide where on the parabola of f(x) you are placing the on-ramp locations. Write those ordered pairs down. 3. West Mathington wants to connect these on-ramps with some surface roads. Create a linear growth function, h(x), that passes through both on-ramp points. Create an exponential growth function, j(x), that passes through at least one of the on-ramp points. Show all of the work you did to create both functions. 4. What important relationship do the x-coordinates of the on-ramp location points have with the system of equations formed by the two roads’ functions that are being connected? Provide justification and support for your explanation. 5. The city planner needs to identify the most northern road. Prove which road will eventually go the furthest to the north (positive y-direction). Create tables for your functions using an appropriate domain of five integers. Using the tables and graph, explain to the city planner which road will be the furthest north as the x-values continue to get larger (the road continues to go east). Provide reasoning why. 6.Include your graph that shows the functions that model each of the roads and the on-ramps.

OpenStudy (anonymous):

don't mind the question marks they aren't anything

OpenStudy (jskhupmang):

ok do you know how to create a function with 2 zeros?

OpenStudy (jskhupmang):

choose 2 points you like on the x-axis for your zeros

OpenStudy (anonymous):

i made this, for #1 f(x)=x^2-1 0=x^2-1, so (1,0) and (-1,0)

OpenStudy (jskhupmang):

nope that wouldn't work because it doesn't have zeros on the x axis

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Can you help me make one or can i adjust that one a little?

OpenStudy (jskhupmang):

nevermind it does work

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

so #1 is done right

OpenStudy (jskhupmang):

yup

OpenStudy (anonymous):

Ok on to #2

OpenStudy (anonymous):

This is what I did

OpenStudy (anonymous):

f(3)= 3^2-1 = 8 (3,8) f(4)= 4^2-1 = 15 (4,15)

OpenStudy (jskhupmang):

ok

OpenStudy (anonymous):

so that works?

OpenStudy (jskhupmang):

ya but its gonna be up there

OpenStudy (anonymous):

oh ok should we bring down the numbers to something else 1 and 2 maybe? or 2 and 3?

OpenStudy (jskhupmang):

that would be nice

OpenStudy (anonymous):

which one?

OpenStudy (jskhupmang):

I would chose like x=-1 and x=2 depends how the freeway you want

OpenStudy (anonymous):

Ok i will do that.

OpenStudy (jskhupmang):

since the zeros were the surface

OpenStudy (anonymous):

i get 0 for f(-1)

OpenStudy (jskhupmang):

yes

OpenStudy (anonymous):

ok and 3 for f(-2)

OpenStudy (jskhupmang):

yup

OpenStudy (anonymous):

ok

OpenStudy (jskhupmang):

So write those ordered pairs down

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

(-1, 0) and (2,3)

OpenStudy (anonymous):

Ok so that is #2 onto number 3

OpenStudy (phi):

h(x) is the equation of a line that passes through your 2 points

OpenStudy (anonymous):

Where'd u go :(

OpenStudy (anonymous):

Ok how about, h(x) = 2x+7

OpenStudy (phi):

which points are you using ?

OpenStudy (anonymous):

hmm (-1,0) and (2,3)

OpenStudy (anonymous):

look above please so you can see where we are at it will help a lot

OpenStudy (anonymous):

thx for staying and helping

OpenStudy (phi):

I would find the slope between those two points: change in y divided by change in x

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

is it 3?

OpenStudy (anonymous):

3-0 ____ 2-1

OpenStudy (phi):

almost. the bottom is 2 - (-1) right ?

OpenStudy (anonymous):

Oh yes my bad the two minuses

OpenStudy (anonymous):

so its 1 3 __ 3

OpenStudy (phi):

yes, so the equation is y = 1x + b or just y= x+b now replace x and y with one of the points (-1,0) for example and find b

OpenStudy (anonymous):

hmmm ok

OpenStudy (anonymous):

-1=0+b

OpenStudy (anonymous):

b=-1

OpenStudy (phi):

I think x=-1 and y is 0 in (-1,0) you did it the other way round.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

ok 0=-1+b

OpenStudy (anonymous):

b=1

OpenStudy (anonymous):

@phi

OpenStudy (phi):

so the equation is y= x+1 let's test: (-1,0) should be "on the line" y= -1+1 = 0 which is correct now check (2,3) y= 2+1= 3, and that is good.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So #3 is done right?

OpenStudy (phi):

you found h(x)= x+1 but wait, there is more: Create an exponential growth function, j(x), that passes through at least one of the on-ramp points.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

j(x)=2x^2+4

OpenStudy (anonymous):

actually no that won't work

OpenStudy (phi):

which point do you want to go through?

OpenStudy (anonymous):

can i go through (2,3)

OpenStudy (phi):

ok remember exponential growth function means base^x I would use 2 for the base (it's simple) so y= 2^x + B we replace x and y with 2 and 3, and find B

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

3=2^2+b b=-1

OpenStudy (phi):

so what equation do you get?

OpenStudy (anonymous):

3=2x^2-1?

OpenStudy (phi):

first, the equation has a y (or in this case , j(x) ) not a 3 also, exponential means x is an exponent. x^2 is a parabola or quadratic which is different. can you try again ?

OpenStudy (anonymous):

ok

OpenStudy (phi):

use the equation I posted up above, but replace B with -1

OpenStudy (anonymous):

j(x) =2^2-1?

OpenStudy (phi):

closer, but we need x as the exponent not 2 (we used x=2 so we could find B, but now we leave it x)

OpenStudy (anonymous):

ok j(x)=2^x-1

OpenStudy (phi):

yes. this equation has an x that we "replace" with a number to find the y for example, when x is 2 we get j(2) = 2^2-1 = 4-1= 3 (as we should) but we can get other points: when x is 0 we get j(0) = 2^0 - 1 = 1-1 = 0 and so on...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok i understand that :)

OpenStudy (anonymous):

@phi still here? :(

OpenStudy (phi):

yes.

OpenStudy (anonymous):

ok thx for sticking with me

OpenStudy (phi):

You can use geogebra to plot your equations

OpenStudy (anonymous):

ok is that for #3?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

so #4 this is where i have no idea what to do.

OpenStudy (anonymous):

4. What important relationship do the x-coordinates of the on-ramp location points have with the system of equations formed by the two roads’ functions that are being connected? Provide justification and support for your explanation.

OpenStudy (phi):

can you see the plot I posted. the x coordinate of the point A (-1,0) is the solution to the system of equations: parabola and line the x value of point B is the solution to the system with parabola, line and exponential

OpenStudy (anonymous):

yes i can

OpenStudy (anonymous):

so that is the answer to 4 right what u wrote

OpenStudy (phi):

yes, but you can write in the equations (not just the names)

OpenStudy (anonymous):

can you explain or write it like that for me i don't know what you mean